

********************************************************************
UARS WEEKLY STATUS REPORT - 19 NOV 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible. We remained on and collecting normal data during the
Leonids event this week.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar data this past week without
apparent problems. Occultation scans still remain suspended. These will
resume if telemetry continues to be stable.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
During this week the MPG has been involved with the evaluation
of the potential impact on instrument operations in case both
tape recorders become unuseable. Without tape recorders, science
data can continue to be collected during real time contacts, but
the expanded time of transponder operations would draw some power
which would otherwise be available to the instruments.
***********
CDHF
***********
The CDHF has notified PI Software Representatives that no Data Gaps were
among the ingested times for UARS DAYS 2983 - 2986. Production processing
has resumed with normal processing results.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
17 days CLA (7.0) PROD (L1-L3) (476R-391C-19F=66L)
22 days GPI (4.05) PROD (CLAES L3B) (334R-307C-2F=25L)
21 days WIN (5.11A) PROD (LOG) (21R-21C=0L)
3 days WIN (5.11A) PROD (L1-L3) (21R-3C-2F=16L)
Promoted WINDII V5.11A software to Production Mode.
DAAC DATA:
- All of the SUSIM Level 3BS (Output 20) UARS DAYS 31-2523 that SUSIM
requested to be sent to the DAAC has been loaded into the DAAC Transfer
Tables. SUSIM and DAAC have been notifed.
CDHF REHOST:
Testing SUSIM V7.1 on the Rehost System.
Preparing to test SOLSTICE V4.62 on the Rehost System.
***********
FOT
***********
REPORTING PERIOD
This report covers 6 November, 1999 (Orbit 44556, GMT Day 310) through 12
November 1999 (Orbit 44660, GMT Day 316).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 314/1352 44624 Off
PEM 314/1954 44627 AXIS On
PEM 315/2322 44645 AXIS-2 Off
The beta angle ranged between the angles of 30.3 degrees to 60.8 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
76 degrees on 18 November 1999 (DOY 322).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.1 degrees C. Additional battery performance data for each
battery on 05 November 1999 and 12 November 1999 are provided in Appendix D.
The clock error ranged from 10.4 to -10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 03
% PEM = 02
% HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/284 - 2287.496981
% Transponder B: 99/265 - 2287.498700
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/315 show the observatory data loss to be 150 hours, 41
minutes, 19.097 seconds (an increase of 66 hours 45 minutes 27.424 seconds
since the last report period). This is a 0.2113 percent data loss which equals
a 99.7887 percent data capture for the mission.
The increase of 66 hours 45 minutes 27.424 seconds occurred over seven (7)
days:
7 05 hours 41 minutes 16.064 seconds on Day 309 due to a S/C TR problem.
7 15 hours 24 minutes 12.608 seconds on Day 310 due to a S/C TR problem.
7 19 hours 36 minutes 28.160 seconds on Day 311 due to a S/C TR problem.
7 19 hours 18 minutes 03.968 seconds on Day 312 due to a S/C TR problem.
7 05 hours 58 minutes 43.168 seconds on Day 313 due to a S/C TR problem.
7 00 hours 46 minutes 19.840 seconds on Day 314 due to a S/C TR problem.
7 00 hours 00 minutes 23.616 seconds on Day 315 due to a S/C TR problem.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Investigation and testing continued with NBTR-B this week. The Spooling
test was in progress since Tues., Nov 2. Tape recorder performance was nominal
until Fri., Nov 5. That evening NBTR-B resumed the same failure pattern and
the FOT was notified. The FOT came in on Sat, Nov 7 and executed the 4-Mux
Mode test. The 4-Mux Mode works in the manner that four of record/playback
heads record in the forward direction, called Pass 1 while the other 4 heads
record in the reverse direction, called Pass 2. If a particular head is
identified to be anomalous, the data from either Pass 1 or 2 can be isolated
and used.
The 4-Mux Mode test on Sat, Nov 7 was NOT entirely successful. Only the data
from Pass 2 was captured. The 4-Mux Mode test was attempted two more times on
Mon., Nov 8 with the same results, Pass 2 data only captured. This indicated
that one or more heads on Pass 1 are intermittent.
The test was continued again on Tues., Nov 9 with data captured from both Pass
1 and 2 with no dump errors observed. After a day of successful
record/playback runs, NBTR-B was left in the 4-Mux Mode and scheduled to
maximize data capture. To date, both pass 1 and 2 data have been captured with
no anomalies to report. The operation is still considered in the test mode,
therefore TDRSS support will continue to be maximized and the anomaly
investigation will continue.
The anomaly investigation continues with data analysis, test plans and the
operational capability. Additional contractual support has been provided to
help this anomaly. Specifically;
- Data analysis has been hindered by the nature of the anomaly. During the
anomaly, the frame sync bits don't matched the expected pattern and the
entire frame is thrown away. This translates to no data capture. Without
data, all of the NBTR-B parameters cannot be analyzed. Currently an action
is in place to release the frame sync bit pattern restrictions and process
all the data, good or bad.
- The Test Plans are continually discussed with their value, content and
hazards reviewed.
- An issue under investigation surrounding the operational capability lies
within the Erase function. According to the provided operating documentation
the only way to erase is to place NBTR-B into record. The Erase function is
automatically enabled during record but only during the forward direction, or
Pass 1. When the recorder switches to Pass 2, Erase is automatically
disabled. What this all translates to is that the recorder must be in record
for four hours to capture two hours of data in a Pass 1 failure scenario. If
Pass 1 becomes intermittent, an Erase function could significantly increase
recorder coverage by eliminating the slow erasure process of recording on all
of Pass 1 just to erase Pass 2. For example, Pass 2 may be able to be erased
in the Wind Mode which is significantly faster than Rec Mode.
- Contractual Support has been appreciated by the investigating FOT. In
general, outside support has aided the FOT by providing insight into the
nature of the anomaly, test plan review, a name database for additional
contacts and help, data analysis, failure history and capabilities of other
missions, documentation and much more.
- On Wed, Nov 10, 1999 the RF switch was configured from RHCP to LHCP on the
transponder at 1500 UTC, from the stored command load. This gives the
capability to take MA supports as well as SSA. The MA service can now be
scheduled to help fill scheduling gaps and maximize realtime. In the event
NBTR-B fails completely, SSA/MA coverage will be the only means for capturing
the science data. The RF switch was previously configured to RHCP to reduce
mutual interference from other mission using LHCP. Over the past few years the
mutual interference has diminished. Since the reconfiguration, no interference
has been reported.
The FOT continued it's effort to schedule additional TDRS contact time in
support of the NBTR B anomaly. This effort involved extending existing SSA
events, as well as adding new SSA and MA events, in an attempt to maximize
real-time communications with the satellite. Due to the polarization switch of
transponder B from Right Hand Circular Polar (RHCP) to Left Hand Circular Polar
(LHCP), all new and rescheduled events (two weeks worth) had to be added using
LHCP configuration codes. This all made for a very intensive week of TDRS
scheduling.
YEAR 2000 OPERABILITY SUPPORT
All areas identified to or by the Flight Operations Team for Y2K testing have
been completed. There are no other areas identified at this time that have not
been Y2K certified.
MEETINGS
DOY 312, Mon. 10 am, Nov 8, 1999 - A briefing was held on the status of NBTR B.
The operating performance and anomaly trends were discussed. These discussions
lead to the identification of contractual engineers with relevant experience.
Permission was given to contact and update the identified support.
DOY 312, Mon. 2 pm, Nov 8, 1999 - A meeting was held to discuss the status of
NBTR-B as well as brief Contracted guests to the events of the past few weeks.
The FOT's test recommendations were discussed. All were in agreement to the
tests that the FOT had recommended and the order and condition in which they
should be completed. Also agreed upon, was the throwing of the RF switch for
MA supports and LHCP SSA events.
DOY 313, Tues. 10:30 am, Nov 9, 1999 - A meeting was held to discuss some of
the details to throw the RF switch. An implementation schedule for the
throwing of the RF switch, LHCP for SSA and MA support with GCMR commands were
discussed. Additional contract support was present and briefed as to the
status of NBTR-B.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group. It is expected that one
more meeting will be held before Y2k as an update to UARS safehold expectations
for the new century.
The Leonid Shower is expected and planned for November 17 through 18, 1999.
Special operational procedures are in place. The FOT will be present to
oversee the implementation of the plan.
UARS has tested and prepared for the Year 2000 event. Special considerations
have been made in preparation for a non nominal 1999 to 2000 transition.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 November 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
44563 310/1325 99-368 OBC Clk Adj 0.00 to 20.6 msec/day
44591 312/0939 99-369 OBC Clk Adj 20.6 to 0.00 msec/day
44518 314/0502 99-370 OBC Clk Adj 0.00 to 20.6 msec/day
44624 314/1352 99-371 HALOE Power OFF
44627 314/1954 99-329 PEM Power ON AXIS 1 & 2
44645 315/2322 99-372 PEM Power OFF AXIS 2
_________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 November 1999
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON R-LO PTS PB I Monitor 44586
HRPSEPBIMON R-LO PSE PB I Monitor 44586
HATBAT Y-HI HALOE Bolo Temp 44620-22
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 November 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
44563 310/13:28 (06 Nov 99) from 00.0 to 20.6 msec/day
44591 312/09:52 (08 Nov 99) from 20.6 to 00.0 msec/day
44618 314/05:16 (10 Nov 99) from 00.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 November 1999
05 Nov 1999 - GMT Day 309 Beta = 30.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ----- ------- ------
1 -89.6/-106.4 -3.05/-4.94 +0.0/+0.0 16.3 0.000 0.0
2 +28.0/-11.2 +5.47/+3.87 +29.6/-15.2 25.3 1.086 21.2
3 +33.6/-22.4 +4.51/+2.60 +29.2/-15.2 25.1 1.083 21.0
12 Nov 1999 - GMT Day 316 Beta = 60.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- --------- ------- ---------- ----- ------- ------
1 -72.8/-95.2 -3.68/-4.94 +0.0/+0.0 16.3 0.000 0.0
2 +22.4/-11.2 +3.87/+2.60 +24.0/-14.4 26.7 1.030 18.4
3 +22.4/-22.4 +2.92/+1.66 +23.2/-14.8 26.6 1.027 17.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 November 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-080 WINDII 44600 312/2323 N/A WINDII Error Code 03
99-082 FOT 44638 315/1201 21565 Wrong DSID for MA event
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-080 WINDII 44600 312/2323 N/A WINDII Error Code 03
99-082 FOT 44638 315/1201 21565 Wrong DSID for MA event
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
99-081 FOT November 14170 Generic Late Acquisition
SEVNTFW99/46